CHEMWORK Scenario: A certain first-order reaction is 66 percent complete in 65 seconds. Calculate the half-life for this reaction 1/2 Submit Submit Answer Try Another Versiorn 10 item attempts remaining
Answer #1

First order kinetic equation is

k = 2.303/t log (a/a-x)

= 2.303/65 log (100/100-66)

= (0.035) (0.477)

= 0.016

Half life period (t1/2) = 0.693/k = 0.693/ 0.016 = 43.3 s